Assume that the market for Good X is defined as follows: QD = 64 - 16P and QS = 16P - 8. If the government imposes a price floor in this market at $3.00, what will consumer surplus be?
Answer #1

Answer

Consumer Surplus is the area below demand curve and above price line.

Price floor will effect market price if price floor is set above equilibrium price.

Equilibrium occurs when QD = QS => 64 - 16P = 16P - 8 => P = 2.25

Hence Price floor = $3 > Equilibrium price, thus market price will be price equal to price floor.

Thus After price price market price = $3

When Price = 3 , QD = 16

When QD = 0 then P = 4(Vertical intercept)

Hence Consumer Surplus = (1/2)(4 - 1)*16 = 24.

Hence, Consumer Surplus will be $24.